OpenStudy (mathmath333):

Rajiv lend out rs 9 to Anni on condition that the amount is payable in 10 months by 10 equal installments of rs. 1 each payable at the start of every month.what is the rate of interest per annum , if the first installment has to be paid one month from the date the loan is availed?

OpenStudy (mathmath333):

books answer \(\Large 2.222\%\)

ganeshie8 (ganeshie8):

it should be way more than 10% right ? because 1 is more than 10% of 9

OpenStudy (mathmath333):

http://prntscr.com/5myr2w

OpenStudy (mathmath333):

i cant understand the book's answer

ganeshie8 (ganeshie8):

I don't get the textbook solution either. Here is my attempt : Suppose the monthly interest rate = \(i\) Balance after \(1\) month = \(9(1+i) - 1\) Balance after \(2\) months = \([9(1+i) - 1](1+i)-1 = 9(1+i)^2 - (1+i)-1\) \(\cdots\) Balance after \(10\) months = \( 9(1+i)^{10} - (1+i)^9-\cdots -1\)

ganeshie8 (ganeshie8):

we want the balance after 10 months to be 0 as the loan gets completely paid off by then : \( 9(1+i)^{10} - (1+i)^9-\cdots -1 = 0\) \( 9(1+i)^{10} = (1+i)^9+\cdots +1 \) right hand side is a geometric series, so we can use the partial sum formula

ganeshie8 (ganeshie8):

for the right hand side, there are 10 terms, first term = 1 and common ratio = 1+i so we get \(9(1+i)^{10} = \dfrac{(1+i)^{10} - 1}{(1+i)-1}\) \(9(1+i)^{10} = \dfrac{(1+i)^{10} - 1}{i}\) we can solve \(i\) i guess

ganeshie8 (ganeshie8):

multiply \(i\) by 12 for the annual interest rate \(r = 12i = 12(0.01963) = 0.23556 = 23.556\%\)

ganeshie8 (ganeshie8):

answer should be 23.556% 2.222% doesnt make sense to me :O

OpenStudy (mathmath333):

oh ok, thnx